OpenStudy (anonymous):

Determine the equation of the tangent to the graph of each curve at the point that corresponds to each value of x. a) y = (6x-3)(-x^2+2) I can't believe Im not getting this right. I don't know what I am doing wrong below: y = -6x^3 + 12x +3x^2 - 6 y' = -18x^2 + 12 + 6x f(1) = 3 f'(1) = 0 y - 3 = 0(x - 1) y = 3. The answer at the back says: y = 3x-11

OpenStudy (amistre64):

the equation for the tangent is: y = m(x-Px)+Py

OpenStudy (amistre64):

tangent = (f'(Px))(x-Px)+Py ; where P(x,y) is the point given

OpenStudy (anonymous):

So what did I do wrong there?

OpenStudy (amistre64):

f(x) = (6x-3)(-x^2+2) f'(x) = 6(-x^2+2) + (6x-3)(-2x) f'(x) = -6x^2 +12 -12x^2 +6x f'(x) = -18x^2 +6x +12 right?

OpenStudy (amistre64):

the point (1,3) is given i think

OpenStudy (anonymous):

when subbed in x = 1 for y', you get 0.

OpenStudy (amistre64):

right, f'(1)=0; which means that the equaion is just y = 3 but are you sure you are looking at the right problem?

OpenStudy (amistre64):

my simple mistakes tend to be becasue I left out a part, forgot a sign, or just wrote down the wrong problemm